how do u graph (x-3)^2 divided by 25 + (y+6)^2 divided by 49 = 1
Answer by lwsshak3(11628)

how do u graph (x-3)^2 divided by 25 + (y+6)^2 divided by 49 = 1
..
Standard form for ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1 (a>b), with (h,k) being the (x,y) coordinates of the center
Standard form for ellipse with vertical major axis: (x-h)^2/b^2+(y-k)^2/a^2=1 (a>b),with (h,k) being the (x,y) coordinates of the center
The only difference between the two forms is that a^2 and b^2 are interchanged
..
Given equation in standard form: (x-3)^2/25+(y+6)^2/49=1
This is an ellipse with center at (3,-6). It has a vertical major axis because the larger denominator is under the y-term. (2nd form listed above)
Other information needed for graphing:
a^2=49
a=7
length of major axis =2a=14
b^2=25
b=5
length of minor axis =2b=10
c^2=a^2-b^2=49-25=24
c=√24=4.9
End points of major axis or location of vertices: (3,-6±7)
Location of foci: (3,-6±√24)
You now have the information you need to graph equation of given ellipse.
Your graph should look much like the graph below:

..
y=(49(1-(x-3)^2/25))^.5-6